OpenStudy (anonymous):

Can someone pls tell me whether the graphs of the lines will be collinear, parallel, or intersecting? y= 4x+1 y=4x-3

jimthompson5910 (jim_thompson5910):

What are the slopes of each line?

OpenStudy (anonymous):

intersecting

OpenStudy (anonymous):

why?

jimthompson5910 (jim_thompson5910):

no, the lines aren't intersecting

OpenStudy (anonymous):

I just need the explaination of the formula to tell whether or not it is parallel or intersecting and col linear

jimthompson5910 (jim_thompson5910):

What are the slopes of each line?

OpenStudy (anonymous):

thats all it is given, theres a formula for solving but I don't know it.

jimthompson5910 (jim_thompson5910):

Hint: the equation y = mx+b has the slope m

jimthompson5910 (jim_thompson5910):

ex: y = 9x + 20 has the slope of 9

OpenStudy (anonymous):

sorry the lines are not intersecting. Extremely sorry, I made a mistake. They have same slopes, so they should be parallel.

OpenStudy (anonymous):

I guess its four right? thats fine

jimthompson5910 (jim_thompson5910):

4 is the slope, yes

jimthompson5910 (jim_thompson5910):

since the slope is 4 for both of the lines, the lines are either parallel or collinear

jimthompson5910 (jim_thompson5910):

the two lines are collinear if and only if the two y-intercepts are the same

jimthompson5910 (jim_thompson5910):

otherwise, the two lines are parallel

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

one passes from y=-3 and other from y=1, so they are parallel(since slopes are same) but not co-linear.

OpenStudy (anonymous):

so if its the same its parallel and what happens if its collinear?

jimthompson5910 (jim_thompson5910):

no, if the y-intercepts are the same, the lines are collinear if the y-intercepts are different, the lines are parallel both cases are true when the slopes are both equal

OpenStudy (anonymous):

oh ok i see, so it is collinear right?

jimthompson5910 (jim_thompson5910):

what are the y-intercepts for both lines?

OpenStudy (anonymous):

i need help finding the y intercepts pls

jimthompson5910 (jim_thompson5910):

y = mx+b b is the y-intercept ------------------------------ ex: for y = 2x+11 11 is the y-intercept ------------------------------ ex: y = 14x-13 which is the same as y = 14x + (-13) -13 is the y-intercept

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

in this case the y intercepts are the same, which is 4x? And what do I do with the 1 and -3?

OpenStudy (anonymous):

to find Y=intercept, put x=0 in each equation. simple. Any point on y-axis is represented as x=0 and y=Y

jimthompson5910 (jim_thompson5910):

no, that's not the y-intercept

jimthompson5910 (jim_thompson5910):

y = 4x + 1, y intercept is 1

jimthompson5910 (jim_thompson5910):

look back up at my 2nd to last post

OpenStudy (anonymous):

oh right then the y intercepts aren't the same

jimthompson5910 (jim_thompson5910):

so the lines are parallel

OpenStudy (anonymous):

ah ok what will happen if its intersecting?

jimthompson5910 (jim_thompson5910):

if the slopes are different, then the lines will intersect

OpenStudy (anonymous):

what about collinear?

jimthompson5910 (jim_thompson5910):

if the slopes are the same and the y-intercepts are the same, then the two lines are collinear

OpenStudy (anonymous):

ah coolio thanks